OpenStudy (anonymous):

Solve the equation on the interval 0≤θ<2π: 2cos²θ-1 = 0

OpenStudy (anonymous):

start with \[\cos^2(x)=\frac{1}{2}\]

OpenStudy (anonymous):

take the square root, don't forget the \(\pm\)

OpenStudy (anonymous):

you get \[\cos(x)=\pm\frac{1}{\sqrt{2}}\] which you might recognize better as \[\cos(x)=\pm\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

then see if you know some angles \(x\) where \[\cos(x)=\frac{\sqrt{2}}{2}\] and also where \[\cos(x)=-\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw you good from there?